Session: May 2002 (session 2)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Xinyong Huang at Thu Jun 6 20:07:34 CEST 2002) Success

Medicina (vargiu at Thu Jun 6 20:27:59 CEST 2002) Success

Urumqi (Chen Maozheng at Fri Jun 7 07:10:46 CEST 2002) Success

Effelsberg (A. Kraus at Fri Jun 7 10:34:43 CEST 2002) Minor failures
A few times, VC02 had an overflow.

Noto (P. Leto at Fri Jun 7 10:50:57 CEST 2002) Success
No known problems

Onsala (John Conway at Sat Jun 8 16:41:23 CEST 2002) Minor failures
At start some instablities in phase cal signal, but was stable for most of expt

Onsala (John Conway at Sat Jun 8 16:43:09 CEST 2002) Minor failures
For unknown reasons the first tape wound off the spool 1 hour before end. Otherwise OK.

\(Other\) (Algonquin at Mon Jun 10 15:12:38 CEST 2002) Success
Algonquin completed ec017a experiment with success,no missed source. Corrected fmout-gps= +1.1095 microseconds Wx temp=15.6c pressure=989.7mb humidity=37.5% sky coverage=30% Tape drive temp=110.7f humidity=23% SEFD 3c123 az=230 el=68 x=190 s=206 sunny,thin clouds 176 199 Receiver front=27.65 20k=31.19 70k=76.13

Hartebeesthoek (Jonathan Quick at Thu Jun 20 10:18:44 CEST 2002) Success
No known problems.